what is volume of a 0.62 mol sample of gas at 10.6 and a pressure of 1.05


1
Expert's answer
2022-03-12T01:59:28-0500

Solution.

"n=0.62 mol;"

"T=10.6^oC=283.6K;"

"P=1.05 atm;"

"V-?;"

"PV=nRT\implies V=\dfrac{nRT}{P};"

"V=\dfrac{0.62 mol\sdot 0.0821 L/atm/mol/K\sdot 283.6K}{1.05 atm}=13.75L;"